40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2918     Entry time: Wed May 12 03:56:54 2010
Author: Koji 
Type: Update 
Category: IOO 
Subject: Faraday aligned 

Zach and Koji

The old small MMT was removed and wrapped by Al foils.

The steering mirror IM2-IM4 were displaced and aligned.

The Faraday isolator block is moved and aligned.

The MC is realigned and resonatng TEM-00.

Now the MC has slightly miscentered beam on the mirrors owing to change of the stack leveling.
OSEMs are also in a strange state. We should check this later.

ELOG V3.1.3-